future value of ordinary annuity


bab savings and loan is paying 6% interest compounded monthly. How much will 1000 deposited at the end of each month be worth after 2 years?


1
Expert's answer
2021-03-24T09:09:36-0400

Since payments of an ordinary annuity are made at the end of a period and the last payment does not earn interest;


"FVOA\ =A\ *\ \frac{{(1+r)}^n-1}{r}"


"rate\ compounded\ monthly\ =\ \frac{0.06}{12}\ =\ 0.005\ interest\ p.m"



"FVOA\ =300,000\ *\ \frac{{(1+0.005)}^{24}-1}{0.005}\ =\ 25,431.96"


"The\ future\ value\ of\ ordinary\ annuity \ will\ be \ \$25,431.96"